Find the vertices of a hyperbola with the equation (x + 3)2 – 4(y – 2)2 = 4.
Answer by lwsshak3(11628)

Find the vertices of a hyperbola with the equation (x + 3)2 – 4(y – 2)2 = 4.
**
(x + 3)^2 – 4(y – 2)^2 = 4
Divide by 4
(x+3)^2/4-(y-2)^2/1=1
This equation is a hyperbola with a horizontal major axis of standard form:
(x-h)^2/a^2-(y-k)^2/b^2=1
For given equation:
center: (-3,2)
a^2=4
a=2
Vertices: (-3ħa,2)
(-3ħ2,2) or (-5,2) and (-1,2)